In an A.P sum of m terms is equal to n and sum of n terms equal to m , then sum of (m+n) terms is

Grade:10

1 Answers

Arun
6 years ago
Sm= n & Sn= m (given)
now m/2(2a+ (m–1)d)= n
2am+ m(m-1)d= 2n     (1)
similarily
2an+ n(n-1)d= 2m       (2)
subtracting (2) from (1)-
we get 2a + (m+n –1)d= –2
now Sm+n= (m+n)/2 [2a+ (m+n-1)d]
Sm+n= – (m+n)   (Answer)
note- please learn this formula as some times this will be used directly in questions. hope this helps